OpenStudy (anonymous):

4 questions? HELP PLEASE 1.) 1/10 (5+5) +3 =? 2.)12/30 (5) (6) + 3=? 3.) .18/.36= ? (leave this as a fraction) 4.) .8/.40=? @ZeHanz

OpenStudy (anonymous):

do you know how to do it

OpenStudy (anonymous):

No :(

OpenStudy (zehanz):

1. First, calculate the part between the brackets, then multiply with 1/10, finally add 3.

OpenStudy (anonymous):

how do i multiply 10 to 1/10?

OpenStudy (zehanz):

suppose you have a cake sliced in 10 pieces. Then every piece is 1/10 of the cake. What do you get when you are offered 10 pieces?

OpenStudy (anonymous):

a whole?

OpenStudy (anonymous):

ohhhhhhhhhhhhhhhhhh

OpenStudy (anonymous):

Mind=blown

OpenStudy (anonymous):

now can you explain it with the decimal :(

OpenStudy (zehanz):

#3? Seems difficult until you realize the denominator is twice the numerator...

OpenStudy (anonymous):

Huh?

OpenStudy (zehanz):

Look at it: numerator: .18 denominator: .36. That is twice as much. Here are some examples: 3/6, 100/200, 2/4, 7/14 Everytime this really is the same number! It is also the same number when you have .18/.36 Do you see what number it is?

OpenStudy (anonymous):

so it would be 2?

OpenStudy (zehanz):

You wouldn say 3/6 is equal to 2, now would you?

OpenStudy (anonymous):

Im so confused *-*

OpenStudy (anonymous):

\[\frac{1}{100}=.01\] \[\frac{1}{10}=.1\] using these you can make up what each decimal is actually in fractions

OpenStudy (anonymous):

so for .18 you have \[\frac{1}{10}+\frac{8}{100}\]

OpenStudy (zehanz):

If you had 3 parts out of 6 how much would you have?

OpenStudy (anonymous):

1/2

OpenStudy (zehanz):

Yes! so just like the other fractions I mentioned, .18/.36 is also 1/2

OpenStudy (anonymous):

so im stuck on number 2

OpenStudy (anonymous):

i dont understand how to multiply 12/30 by 30

OpenStudy (zehanz):

Sorry, going to sleep now... first do 5*6 then 12/30 * that outcome then+3

OpenStudy (anonymous):

mhmmk bye

OpenStudy (anonymous):

is 2 this \[\frac{12}{30}*(5)*(6)+3\]

OpenStudy (anonymous):

yes

OpenStudy (zehanz):

OK last one: see it another way: youhave 12/30 * 30 So you have 12, then you divide by 30, then you multply with 30. What do you get ?

OpenStudy (anonymous):

divide 12 by 30?

OpenStudy (zehanz):

Yes and directly after that *30...

OpenStudy (anonymous):

12?

OpenStudy (zehanz):

Yes!, the operations cancel!

OpenStudy (anonymous):

look at is as this \[\frac{12}{30}*\frac{30}{1}\] this is the same as \[\frac{12}{1}*\frac{30}{30}\]... \[\frac{30}{30}=1\]

OpenStudy (zehanz):

Bye

OpenStudy (anonymous):

so you're left with \[12\]

OpenStudy (anonymous):

so the answer would be 12. Mhmmk

OpenStudy (anonymous):

yep in other words if you are multiplying two things and you have the same number in the numerator and denominator you can cancel them since they will equal 1

OpenStudy (anonymous):

then for the last one? .8/.40=? how do i do that

OpenStudy (anonymous):